Determine whether the following functions are continuous on the given interval. Show your complete solution.

1. Ζ’(π‘₯)=√4βˆ’π‘₯2 ;[βˆ’2,2]

2. Ζ’(π‘₯)=3π‘₯2 βˆ’π‘₯+5 ; (βˆ’βˆž,+∞)


1
Expert's answer
2022-03-21T15:41:45-0400

"f\left( x \right) =\sqrt{4-x^2}" is continuous on "D\left( f \right) =\left[ -2,2 \right]" , since it is a composition of polynomial function "4-x^2" and power function "\sqrt{x}"

"f\left( x \right) =3x^2-x+5" is continuous on "\left( -\infty ,+\infty \right)" since it is polynomial function.